# Probability-Biased Attention over Directed Bipartite Graphs for Long-Tail ICD Coding

### v1

@@ Abstract
Automated International Classification of Diseases (ICD) coding aims to assign multiple disease codes to clinical documents, constituting a crucial multi-label text classification task in healthcare informatics. However, the task is challenging due to its large label space (10,000 to 20,000 codes) and long-tail distribution, where a few codes dominate while many rare codes lack sufficient training data. To address this, we propose a learning method that models fine-grained co-occurrence relationships among codes. Specifically, we construct a Directed Bipartite Graph Encoder with disjoint sets of common and rare code nodes. To facilitate a one-way information flow, edges are directed exclusively from common to rare codes. The nature of these connections is defined by a probability-based bias, which is derived from the conditional probability of a common code co-occurring given the presence of a rare code. This bias is then injected into the encoder’s attention module, a process we term Co-occurrence Encoding. This structure empowers the graph encoder to enrich rare code representations by aggregating latent comorbidity information reflected in the statistical co-occurrence of their common counterparts. To ensure high-quality input to the graph, we utilize a large language model (LLM) to generate comprehensive descriptions for codes, enriching initial embeddings with clinical context and comorbidity information, serving as external knowledge for the statistical co-occurrence relationships in the code system. Experiments on three automated ICD coding benchmark datasets demonstrate that our method achieves state-of-the-art performance with particularly notable improvements in Macro-F1, which is the key metric for long-tail classification.
@@
ICD Coding , Multi-label Text Classification , Long-tail , Graph Transformer , Large Language Models
@@ 1 Introduction
The International Classification of Diseases (ICD [^1] https://www.who.int/standards/classifications/classification-of-diseases) is a healthcare classification system maintained by the World Health Organization (WHO) that provides a unique code for each diagnose or procedure. Although widely adopted, the manual process of assigning ICD codes to clinical text is time-consuming, error-prone, and expensive. Automated ICD coding is typically framed as a multi-label text classification task [1], but it presents a significant challenge: a small subset of codes, such as those for chronic conditions like diabetes, dominates the frequency distribution. Meanwhile, the majority of codes, often corresponding to rare diseases or nuanced complications, appear infrequently, creating a long-tail distribution. In the MIMIC-IV-ICD-10 dataset, for example, over 70% of its 26,096 distinct codes occur fewer than 10 times [2]. This severe data imbalance hinders the ability of data-driven classification models to learn meaningful semantic representations for these rare codes.
A primary strategy for automated ICD coding is leveraging co-occurrence relationship (see Figure 1) to capture complicated correlations among codes. Cao et al. [3], Luo et al. [4], and Wang et al. [5] model these correlations via graph networks, motivated by comorbidity as the clinical foundation underlying statistical co-occurrence [6]. However, these methods critically oversimplify co-occurrence relationships into binary connections, ignoring fine-grained statistical information, such as conditional probabilities that refine the correlation between rare and common codes, limiting representation enhancement especially for data-scarce rare codes.
@@
Figure 1: An example of a clinical document assigned the co-occurring codes “Urea cycle disorder", “Acidosis", and others. “Acidosis" is often triggered by “Urea cycle disorder", as toxic ammonia buildup disrupts cellular metabolism.
@@
Another line of research focuses on exploring external knowledge to enrich code semantics. Representative approaches like MSMN [7], KEPT [8] and MSAM [9] expand code representations by aligning code names with synonyms from medical ontologies (e.g., UMLS, Wikidata). While this leverages lexical variations, a critical limitation persists: synonym expansions frequently exhibit semantic homogeneity with the original code names. To illustrate, the synonyms “disorder of urea cycle metabolism, unspecified” and “disorder of the urea cycle metabolism, nos” both belong to ICD code 270.6, which has the name “urea cycle disorder ”. Consequently, such methods are only marginally effective at injecting meaningful knowledge to enhance code representations, leaving significant room for improvements in classification performance.
Additionally, some studies have designed complex multi-stage frameworks. To manage the vast label space of the ICD code system, methods such as GPsoap [10] and TwoStage [11], typically involve a pipeline of discrete steps like retrieving and re-ranking candidate codes. Notably, these frameworks can be combined with other high-performing end-to-end models to achieve better performance.
@@
In this work, we propose a learning method that utilizes fine-grained co-occurrence relationships to enhance rare code representations. Specifically, we construct a Directed Bipartite Graph Encoder with two disjoint node sets: common codes and rare codes. The edges between them are directed exclusively from common to rare codes, with nature identified by conditional probability—the likelihood of a common code occurring given the presence of a specific rare code. We discretize the continuous probability values into bins, which then index a learnable bias in an attention module, a process we term “Co-occurrence Encoding". This graph encoder enhances the representation learning of rare codes by enabling them to aggregate latent comorbidity information from their co-occurring common counterparts. To create high-quality input to the graph, we leverage a specialized LLM prompt to generate comprehensive ICD code descriptions. This approach moves beyond the simple synonyms used in prior work by incorporating detailed clinical contexts and valuable comorbidity information, underpinning the statistical co-occurrence analysis with external knowledge. After encoding code representations, we employ a multi-label attention to get final prediction for each document. We call our model ProBias [^2] The code for our ProBias model is publicly available at https://github.com/ChTianlEI/ProBias-for-ICD-Coding. (derived from the core component “Probability-based Bias" in our proposed Directed Bipartite Graph Encoder). Experiments show that these measures enable our approach to achieve state-of-the-art (SOTA) performance across three benchmark datasets, particularly on metrics sensitive to long-tail distributions.
Our work makes three key contributions:
  • 1.
    First, to the best of our knowledge, this is the first study to enhance rare code representations by directed leveraging information from their co-occurring common counterparts.
  • 2.
    Second, we introduce a novel Directed Bipartite Graph Encoder that integrates a tailored attention mechanism with a probability-based bias. This structure effectively models and utilizes statistical co-occurrence relationship to facilitate information flow from common to rare codes, effectively addressing the long-tail challenge.
  • 3.
    Third, we strategically employ a Large Language Model (LLM) to generate comprehensive descriptions for ICD codes, including clinical context and comorbidity information, effectively augmenting the initial code embeddings with domain-specific knowledge before they are processed by our graph encoder.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work, Section 3 details our methodology, Section 4 presents experiments, and Section 5 concludes with limitations and future directions.
@@ 2.1 Co-occurrence Relationships Extraction
Many existing works utilize co-occurrence among codes to extract their correlations. HyperCore [3] was the first work to exploit code co-occurrence relationships for automatic ICD coding. It defines the co-occurrence using a adjacency matrix, where the edge weights are directly determined by the raw co-occurrence counts between codes, and this structural information is then encoded by a Graph Convolutional Network (GCN). Other approaches have built upon this concept with different strategies for modeling co-occurrence. For instance, Wang et al. [5] propose a method that binarizes co-occurrence relationships by representing the presence or absence of code pairs. These binarized features are then incorporated into a graph transformer model. In another direction, CoRelation [4] integrates the ICD code ontology with the co-occurrence graph to construct a simplified, contextualized graph for each note. To manage complexity, this graph’s nodes are reduced by only selecting the top-K codes with the highest initial prediction probabilities for one node set, while using coarse-grained major code categories for the other.
2.2 External Knowledge Exploration
To enhance the semantic representation of ICD codes, a primary strategy involves extending code names with information from external medical knowledge bases. One prominent approach is to extract synonyms. For instance, MSMN [7] aligns ICD codes with concepts in UMLS to gather synonyms and proposes a multi-synonym attention mechanism where each synonym acts as a separate query to extract relevant text snippets from the clinical note. Further extending this idea, MSAM [9] expands the knowledge sources to include Wikidata and Wikipedia. Recognizing that synonym lists can be repetitive, it also employs a selection algorithm to choose a more representative subset of synonyms for each code. Other works integrate multiple types of knowledge simultaneously. For example, KEPT [8] integrates a pre-trained language model with three domain-specific knowledge sources: code hierarchy, synonyms, and abbreviations.
2.3 Graph Models Utilization
Graph models play a crucial role in automatic ICD coding. For instance, Graph Convolutional Networks (GCNs) [12] have been adopted by ZAGCNN [13], MSATT-KG [14], and HyperCore [3]. More recently, Transformer-based graph models have gained increasing attention. For example, CoRelation [4] employs a graph Transformer [15] to model the interactions between codes on its constructed simplified graph. Similarly, Wang et al. [5] leverage Graphormer [16] to extract the co-occurrence relationships simply based on the connectivity between any two codes.
A key innovation of the Transformer-based graph model Graphormer [16] is how it injects structural information into the attention module. Its “Spatial Encoding" component adds a learnable bias to the attention score between any two nodes. This bias is specifically indexed by the shortest path distance (SPD) between the nodes, allowing the model to incorporate graph topology directly into its calculations. Our work adapts this core principle for the specific challenges of ICD coding.
@@ 3 Method
For automated ICD coding, existing methods commonly treat it as a multi-label text classification problem. Given a clinical document (e.g., discharge summary) $X=\{x_{i}:i\in\{1,2,\cdots,T\}\}$, where $T$ represents the potentially thousands of tokens that compose $X$. Each document $X$ is associated with a ground-truth label set $L=\{l_{i}:i\in\{1,2,\cdots,M\}\}$, where $M$ indicating the number of labels. The objective of automated ICD coding is to predict the label set $L$ for document $X$. Note that, each ICD code represents either a diagnose or a procedure, serving as the task’s label, and typically comes with an original name.
@@
In this section, we introduce the technical details of the proposed method. The pipeline is illustrated in Figure 2.
3.1 Documents Encoding
To start with, we choose GatorTron [17] as the foundational Language Model (LM) for clinical documents encoding. GatorTron is an encoder model pre-trained on medical corpora, which is publicly available in the NVIDIA NGC [^3] https://catalog.ngc.nvidia.com/ Catalog and also through the HuggingFace [^4] https://huggingface.co/UFNLP/gatortron-base library.
As noted in [18], Chunk encoding effectively mitigates the challenges language models encounter when processing lengthy inputs. Given the substantial length of clinical documents, we adopted the approach in [9], partitioning each document into $U$ overlapped chunks and processing each chunk independently using GatorTron. This approach circumvents the token limit of the standard Transformer encoder, enabling effective analysis of lengthy clinical documents. Specifically, for a document $X=\{X_{1},X_{2},...,X_{U}\}$ that has been divided into chunks, where each chunk $X_{u}=\{w_{1}^{u},w_{2}^{u},...,w_{T}^{u}\}$, $U$ has a fixed length of $T=512$ tokens. we employ the GatorTron model to derive its representation:
@@ 3.1 Documents Encoding
3.2 Initial Code Embeddings
To generate the code descriptions, we employ GPT-4o, a brilliant large language model from OpenAI. Its advanced capabilities in semantic understanding, instruction-following, and its extensive knowledge base are crucial for our task, which requires generating highly structured and medically accurate descriptions. To ensure consistency and reproducibility in the output, we set the temperature parameter to a low value (0.2). The model is guided by a structured prompt template designed to produce a comprehensive description for each ICD code, covering clinical contexts, procedural methods, and comorbidity information (see Figure 3 for prompt details and Figure 4 for an example of description).
@@ 3.2 Initial Code Embeddings
Here, $d_{i}$ denotes the description of the $i$-th code $l_{i}$, where $i\in\{1,2,...,N\}$ and $N$ is the total number of codes. These initial code embeddings are then input into the Directed Bipartite Graph Encoder, serving as the basis for capturing fine-grained co-occurrence relationships. Notably, the GatorTron model for these code descriptions remains untrained to mitigate significant GPU memory usage, given the large number of codes.
@@ 3.3.1 Graph Construction
As Figure 2 illustrates, the original code co-occurrence relationships are straightforward: two codes are connected if they co-occur within a document. We calculate a conditional probability metric between rare and common codes:
@@ 3.3.2 Co-occurrence Encoding with Probability-based Bias
To effectively model the co-occurrence relationships between ICD codes, we introduce a novel “Co-occurrence Encoding.” This item is a domain-specific adaptation of the “Spatial Encoding” component in the graph Transformer model Graphormer [16].
While Transformer layers for sequential data explicitly model positional dependencies, often through methods like relative positional encoding [19, 20], graphs lack an inherent sequential arrangement. To address this, Graphormer incorporates a learnable “Spatial Encoding” bias within its attention module, indexed by the shortest path distance (SPD) between nodes. In the medical domain, however, the statistical co-occurrence of codes, which often reflects underlying comorbidity among diseases, serves as a powerful signal [6]. Consequently, our proposed bias is indexed not by distance, but by the conditional probability between code pairs. This allows our model to acquire a more fine-grained, representation-level understanding of co-occurrence relationships among codes.
The implementation of Co-occurrence Encoding involves discretizing continuous probability values into a fixed number of bins. The procedure for generating the bin index $\varphi(l_{r_{i}},l_{c_{j}})$ for any code pair is based on quantile binning (detailed in Algorithm 1).
@@ 3. Compute quantiles: begin
The probability-based bias $c_{\varphi(l_{r_{i}},l_{c_{j}})}$ is a learnable scalar indexed directly by $\varphi(l_{r_{i}},l_{c_{j}})$. This bias is then incorporated into the attention calculation for the $k$-th head ($k\in{1,...,K}$, where $K$ is the number of heads) between a rare $l_{r_{i}}$ code and a common code $l_{c_{j}}$as follows:
@@
The output of the multi-head attention mechanism for the rare code $l_{r_{i}}$ is then the concatenation of the outputs from all heads, followed by a linear transformation:
@@
where $head_{h}=\sum_{j}A^{h}_{l_{r_{i}}l_{c_{j}}}V_{l_{c_{j}}}W^{h}_{V}$, $W_{V}$ is the projection matrix for $l_{c_{j}}$, and $W_{att}$ is the linear layer that combines the outputs of all heads. The resulting rare code attention features, $Att_{l_{r_{i}}}$, along with common code features, $V_{l_{c_{j}}}$, are then processed through a Dropout Layer, Normalization, Residual Connection, and MLP layers to generate the final graph encoder output, $V^{G}=\{V_{i}^{G}\}$, where $i\in\{1,2,...,N\}$ and $N$ is the total number of codes. This feed-forward layer performs local refinement on all codes, which enhances the discriminative power of the output features and improves classification recognition capability.

### v2

# Enhancing Rare Codes via Probability-Biased Directed Graph Attention for Long-Tail ICD Coding

@@ Abstract
Automated international classification of diseases (ICD) coding aims to assign multiple disease codes to clinical documents and plays a critical role in healthcare informatics. However, its performance is hindered by the extreme long-tail distribution of the ICD ontology, where a few common codes dominate while thousands of rare codes have very few examples. To address this issue, we propose a Probability-Biased Directed Graph Attention model (ProBias) that partitions codes into common and rare sets and allows information to flow only from common to rare codes. Edge weights are determined by conditional co-occurrence probabilities, which guide the attention mechanism to enrich rare-code representations with clinically related signals. To provide higher-quality semantic representations as model inputs, we further employ large language models to generate enriched textual descriptions for ICD codes, offering external clinical context that complements statistical co-occurrence signals. Applied to automated ICD coding, our approach significantly improves the representation and prediction of rare codes, achieving state-of-the-art performance on three benchmark datasets. In particular, we observe substantial gains in macro-averaged F1 score, a key metric for long-tail classification.
@@
Graph Attention Networks , Long-Tail Learning , Multi-Label Classification
@@ 1 Introduction
The international classification of diseases (ICD) is a standardized coding system maintained by the World Health Organization (WHO) that assigns unique codes to diagnoses and procedures. Although widely adopted, the manual assignment of ICD codes to clinical documents is labor-intensive, error-prone, and costly, making automated ICD coding an essential tool for modern healthcare systems. Automated ICD coding is typically formulated as a multi-label text classification task [8], because a single clinical document usually corresponds to multiple diagnoses and procedures (see Figure 1 for example), and therefore must be assigned multiple ICD codes simultaneously.
@@
Figure 1: An example of a clinical document assigned with multiple codes “Urea cycle disorder", “Acidosis", and others. “Acidosis" is often triggered by “Urea cycle disorder", as toxic ammonia buildup disrupts cellular metabolism.
Automated ICD coding is particularly challenging due to the extreme imbalance in the ICD label space. A small subset of codes, such as those for common chronic conditions, occurs very frequently, whereas the vast majority correspond to rare diseases and appear infrequently. If we view code frequencies as a distribution, this pattern exhibits a long-tail distribution, where a few common codes dominate the dataset while thousands of rare codes appear only a handful of times. For example, in the MIMIC-IV-ICD-10 dataset, over 70% of the 26,096 distinct codes occur fewer than 10 times [14]. This severe long-tail distribution makes it difficult for data-driven models to learn meaningful semantic representations for rare codes, ultimately limiting overall coding performance.
@@
To tackle the multi-label ICD coding problem, some previous studies adopt multi-stage frameworks to handle the ICD system’s vast label space. Approaches like GPsoap [27] and TwoStage [13] decompose the task into discrete steps, such as candidate code retrieval followed by re-ranking. Another line of research focuses on incorporating external knowledge to enrich code semantics. Methods like MSMN [30], KEPT [28], and MSAM [3] enhance code representations by aligning code names with synonyms from medical ontologies (e.g., Wikidata). While such synonym-based expansion captures lexical variations, it often adds little semantic diversity. For example, the synonyms “disorder of urea cycle metabolism, unspecified” and “disorder of the urea cycle metabolism, nos” both belong to ICD code 270.6 (“urea cycle disorder”). Consequently, such methods are only marginally effective at injecting meaningful knowledge to enhance code representations, leaving significant room for improvements in classification performance.
Another commonly used strategy is to leverage co-occurrence relationships to capture the complex correlations among codes. Several studies model these correlations using graph networks, motivated by comorbidity as the clinical foundation underlying statistical co-occurrence [18, 1, 11, 22]. However, these approaches often oversimplify co-occurrence as binary connections, ignoring fine-grained information such as the actual co-occurrence probabilities. This limitation reduces the effectiveness of representation learning, particularly for rare codes with scarce training data.
@@
In this work, we propose a method that leverages fine-grained co-occurrence information to improve representation learning for rare ICD codes. We construct a directed bipartite graph encoder with two disjoint node sets (i.e., common codes and rare codes), where edges are directed from common to rare codes and weighted by their conditional co-occurrence probabilities. These probabilities are discretized into bins, which index a learnable probability-biased attention term, a mechanism we refer to as co-occurrence encoding. This design allows rare codes to aggregate latent comorbidity information from their frequently observed common counterparts, yielding richer semantic representations. To provide high-quality textual input for the directed bipartite graph encoder, we prompt a medical-domain large language model (LLM) to generate detailed ICD code descriptions that go beyond simple synonyms used in prior work. It can incorporate richer clinical context and comorbidity cues that support more reliable statistical analysis. The resulting code representations are then integrated with a multi-label attention module to produce final predictions. We name our model ProBias [^1] The code for our ProBias model is publicly available at https://github.com/ChTianlEI/ProBias-for-ICD-Coding., derived from its probability-based attention bias. Experiments on three benchmark datasets show that ProBias consistently achieves state-of-the-art performance, delivering substantial improvements on metrics that are particularly sensitive to long-tail label distributions.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 details our methodology. Section 4 presents experiments. Section 5 concludes with limitations and future directions.
@@ 2 Related work
In this section, we review related work from three perspectives: graph-based models, external knowledge integration, and co-occurrence–driven methods.
2.1 Graph-based Models
Graph models play a crucial role in automatic ICD coding. For example, Graph Convolutional Networks (GCNs) [7] have been adopted by ZAGCNN [16], MSATT-KG [25], and HyperCore [1]. More recently, Transformer-based graph models have gained increasing attention. For example, CoRelation [11] employs a graph Transformer [2] to model the interactions between codes on its constructed simplified graph. Similarly, Wang et al. [22] leverage Graphormer [29] to extract the co-occurrence relationships simply based on the connectivity between any two codes. A key innovation of the Transformer-based graph model Graphormer [29] is how it injects structural information into the attention module. Its “Spatial Encoding" component adds a learnable bias to the attention score between any two nodes. This bias is specifically indexed by the shortest path distance (SPD) between the nodes, allowing the model to incorporate graph topology directly into its calculations.
2.2 External Knowledge Integration
To enhance the semantic representation of ICD codes, a primary strategy involves extending code names with information from external medical knowledge bases. One prominent approach is to extract synonyms. For example, MSMN [30] aligns ICD codes with concepts in UMLS to gather synonyms and proposes a multi-synonym attention mechanism where each synonym acts as a separate query to extract relevant text snippets from the clinical note. In addition, MSAM [3] expands the knowledge sources to include Wikidata and Wikipedia. Recognizing that synonym lists can be repetitive, it also employs a selection algorithm to choose a more representative subset of synonyms for each code. Other works integrate multiple types of knowledge simultaneously. For example, KEPT [28] integrates a pre-trained language model with three domain-specific knowledge sources: code hierarchy, synonyms, and abbreviations.
@@ 2.3 Co-occurrence Relationships Extraction
Many existing works utilize co-occurrence among codes to extract their correlations. HyperCore [1] was the first work to exploit code co-occurrence relationships for automatic ICD coding. It defines the co-occurrence using an adjacency matrix, where the edge weights are directly determined by the raw co-occurrence counts between codes. This structural information is then encoded by a Graph Convolutional Network (GCN). Other approaches have built upon this concept with different strategies for modeling co-occurrence. For instance, Wang et al. [22] proposed a method that binarizes co-occurrence relationships by representing the presence or absence of code pairs. These binarized features are then incorporated into a graph transformer model. In another direction, CoRelation [11] integrates the ICD code ontology with the co-occurrence graph to construct a simplified and contextualized graph for each note. To manage complexity, this graph’s nodes are reduced by only selecting the top-K codes with the highest initial prediction probabilities for one node set, while using coarse-grained major code categories for the other.
@@ 3 Method
For automated ICD coding, existing methods commonly treat it as a multi-label text classification problem. Given a clinical document (e.g., discharge summary) $X=\{x_{i}:1\leq i\leq T\}$, where $T$ represents the potentially thousands of tokens that compose $X$. Each document $X$ is associated with a ground-truth label set $L=\{l_{i}:1\leq i\leq M\}$, where $M$ indicating the number of labels. The objective of automated ICD coding is to predict the label set $L$ for document $X$. Note that, each ICD code represents either a diagnose or a procedure, serving as the task’s label, and typically comes with an original name.
In this section, we present the technical details of the proposed ProBias model. The overall pipeline is shown in Figure 2. As shown, the method consists of four main components. First, we generate enriched textual descriptions for each ICD code using LLM, producing high-quality semantic inputs. Second, we construct a directed bipartite graph based on the original co-occurrence statistics and the conditional probability matrix, where edges are directed from common to rare codes and weighted according to different probability levels. Third, we introduce a co-occurrence encoding mechanism that enables information to flow through this graph by injecting a learnable probability-based bias into the attention scores, thereby enhancing the representations of data-scarce rare codes. Finally, we obtain document-level predictions using a co-occurrence-infused multi-label attention module, which aggregates code-aware signals to produce the final ICD assignments.
@@
3.1 Encoding of Documents and Codes
To start with, the clinical documents and ICD codes should be converted into embeddings, with documents encoded via a pre-trained language model and codes represented using LLM-enhanced textual descriptions.
3.1.1 Documents encoding
We choose GatorTron [26] as the foundational language model for clinical documents encoding. GatorTron is an encoder model pre-trained on medical corpora, which is publicly available in the NVIDIA NGC [^2] https://catalog.ngc.nvidia.com/ Catalog and also through the HuggingFace [^3] https://huggingface.co/UFNLP/gatortron-base library.
As noted in [9], Chunk encoding effectively mitigates the challenges language models encounter when processing lengthy inputs. Given the substantial length of clinical documents, we adopt the approach in [3], partitioning each document into $U$ overlapped chunks and processing each chunk independently using GatorTron. This approach circumvents the token limit of the standard Transformer encoder, enabling effective analysis of lengthy clinical documents. Specifically, for a document $X=\{X_{1},X_{2},...,X_{U}\}$ that has been divided into $U$ chunks, where each chunk $X_{u}=\{w_{1}^{u},w_{2}^{u},...,w_{T}^{u}\}$ has a fixed length of $T=512$ tokens. We employ the GatorTron model to derive its representation:
@@ 3.1.1 Documents encoding
3.1.2 Code encoding
For code encoding, we first generate ICD code descriptions using GPT-4o, a state-of-the-art large language model from OpenAI. Its strong semantic understanding, instruction-following ability, and extensive medical knowledge enable the creation of highly structured and clinically accurate descriptions. To ensure consistent and reproducible outputs, we set the temperature parameter to 0.2. The model is guided by a structured prompt template designed to produce comprehensive descriptions for each ICD code, encompassing clinical context, procedural details, and comorbidity information. See Figure 3 for the prompt and Figure 4 for an example.
@@ 3.1.2 Code encoding
Here, $d_{i}$ denotes the description of the $i$-th code $l_{i}$, where $i\in\{1,2,...,N\}$ and $N$ is the total number of codes. These initial code embeddings are then input into the directed bipartite graph encoder (as descriped in the next subsection), serving as the basis for capturing fine-grained co-occurrence relationships. Notably, the GatorTron model for these code descriptions remains untrained to mitigate significant GPU memory usage, given the large number of codes.
@@ 3.2.1 Graph construction
As shown by Figure 2, the original code co-occurrence relationships are straightforward: two codes are connected if they co-occur within a document. We calculate a conditional probability metric between rare and common codes:
@@ 3.2.2 Co-occurrence Encoding with Probability-based Bias
To effectively model the co-occurrence relationships between ICD codes, we introduce a novel co-occurrence encoding. This item is a domain-specific adaptation of the spatial encoding component in the graph transformer model Graphormer [29].
Transformer layers for sequential data explicitly capture positional dependencies, often via techniques like relative positional encoding [17, 15]. However, the graphs lack an inherent sequential order. To address this problem, Graphormer introduces a learnable spatial encoding bias in the attention module, indexed by the shortest path distance (SPD) between nodes. However in the medical domain, the statistical co-occurrence of ICD codes often reflects underlying comorbidity among diseases and provides a more informative signal [18]. Accordingly, our proposed bias is indexed not by distance but by the conditional probability between code pairs, which enables the model to capture a finer-grained, representation-level understanding of co-occurrence relationships. The implementation of co-occurrence encoding is implemented by discretizing continuous probability values into a fixed number of bins. The bin index $\varphi(l_{r_{i}},l_{c_{j}})$ for any code pair is generated via quantile binning, as detailed in Algorithm 1.
@@ 3. Compute quantiles: begin
The probability-based bias $c_{\varphi(l_{r_{i}},l_{c_{j}})}$ is a learnable scalar indexed directly by $\varphi(l_{r_{i}},l_{c_{j}})$. This bias is then incorporated into the attention calculation for the $k$-th head ($k\in{1,...,K}$, where $K$ is the number of heads) between a rare code $l_{r_{i}}$ and a common code $l_{c_{j}}$as follows:
@@
The output of the multi-head attention mechanism for the rare code $l_{r_{i}}$ is obtained by concatenating the outputs from all attention heads, followed by a linear transformation:
@@
Here $head_{h}=\sum_{j}A^{h}_{l_{r_{i}}l_{c_{j}}}V_{l_{c_{j}}}W^{h}_{V}$, $W_{V}$ is the projection matrix for $l_{c_{j}}$, and $W_{att}$ is the linear layer that combines the outputs of all heads. The resulting rare code attention features, $Att_{l_{r_{i}}}$, along with common code features, $V_{l_{c_{j}}}$, are then processed through a dropout layer, normalization, residual connection, and MLP layers to generate the final graph encoder output, $V^{G}=\{V_{i}^{G}\}$, where $i\in\{1,2,...,N\}$ and $N$ is the total number of codes. This feed-forward layer performs local refinement on all codes, which enhances the discriminative power of the output features and improves classification recognition capability.